OpenStudy (anonymous):

Can somebody tell me if I'm correct? 1. Use natural logarithms to solVe the equation. Round to the nearest thousandth. 3e^(2x)+4=24 I got the answer 0.949 2. Write the expression as a single logarithm. 4logb y + 3logb x I got the answer logb(y^4x^3)

OpenStudy (anonymous):

@calculusfunctions

OpenStudy (perl):

3e^(2x)+4=24 3e^(2x)=20 e^(2x) = 20/3 Did you get this far?

OpenStudy (calculusfunctions):

Well first, do you mean\[3e ^{2x +4}=24\]OR\[3e ^{2x}+4=24\]??

OpenStudy (anonymous):

The second one

OpenStudy (calculusfunctions):

OK then first subtract 4 from both sides. Then take the natural logarithm of both sides and remember that\[\ln e =1\]Then solve for x. Also you asked for my help, so if I'm going to tech then you have to please ask @perl and others to stop showing off by just posting the full solutions! If they are actually explaining and teaching, then that's fine.

OpenStudy (calculusfunctions):

OK so now that you have already seen @perl 's intermediary steps, can you finish it by taking the natural logarithm of both sides and solving for x?

OpenStudy (anonymous):

Well I already did the problem and got 0.949. I just wanted to make sure i got the right answer cuz that would mean I did it correct.

OpenStudy (calculusfunctions):

That is the correct answer so I guess you know the steps. Good!

OpenStudy (calculusfunctions):

And yes, you have the answer to next problem also right.

OpenStudy (anonymous):

Decided I would try doing them myself cuz it's my practice test and I should probably take it seriously so I do okay on my test.

OpenStudy (calculusfunctions):

That is a very good attitude!! As a teacher I'm proud! Keep at it...GOOD JOB!!!

OpenStudy (anonymous):

Thanks for the help, math has always been my weak subject im better at English and history. Hoping to be a lawyer. But to do that I have to get through math cuz colleges look at everything.

OpenStudy (calculusfunctions):

Plus it's always good to have at least some background in math and science, regardless of your career choice. The ancient Greeks defined an educated person as someone who had some knowledge of math and science. Good luck to you in whatever you pursue!

OpenStudy (anonymous):

Well I'm in honors math and science and have a B in both.

OpenStudy (anonymous):

Thanks :)

OpenStudy (calculusfunctions):

That's not bad. You're welcome.

OpenStudy (anonymous):

I'm in 10th grade and I have 12th grade Honors English with a 98%. Definitely my strongest subject.

OpenStudy (anonymous):

Last year I had a perfect score on my English final exams.

OpenStudy (calculusfunctions):

WOW!! Good for you!! You're just like my sister then, in that regards.

OpenStudy (calculusfunctions):

Well, if there's nothing else, I'll be logging out for now.

OpenStudy (anonymous):

have a nice night :)

OpenStudy (anonymous):

Or day....it's 3 am here.

OpenStudy (calculusfunctions):

You also.